### v1

# Are heuristic switches necessary to control artificial viscosity in modern smoothed particle hydrodynamics?

Abstract
Context. Artificial viscosity is commonly employed in Smoothed Particle Hydrodynamics (SPH) to model dissipation in hydrodynamic simulations. However, its practical implementation relies on complex numerical switches to restrict its application to regions where dissipation is physically warranted, such as shocks. These switches, while essential, are imperfect and can introduce additional numerical noise.
Aims. Develop and validate a more efficient artificial viscosity scheme for SPH that does not rely on heuristic switches.
Methods. Recent studies have proposed that subtracting the linear component of the velocity field can suppress spurious dissipation in shear-dominated regions. Building on this idea, we implement a velocity-reconstruction technique that removes the bulk linear motion from the local velocity field and uses the Balsara correction to modulate the dissipation.
Results. The presented methodology yields a balanced dissipation scheme that performs well across a range of regimes, including subsonic instabilities, shear flows, and strong shocks. We demonstrate that this approach yields improved accuracy and lower spurious dissipation compared to conventional artificial viscosity switches.
Key Words.:
Methods: numerical: -hydrodynamics – shocks – instabilities
1 Introduction
Smoothed Particle Hydrodynamics (SPH), originally introduced by Gingold & Monaghan (1977) and Lucy (1977), is a Lagrangian method widely used to simulate fluid flows, particularly in astrophysical contexts where complex geometries and free surfaces are common. Despite its versatility and continued development, SPH still faces challenges in certain areas. Notably, the accurate treatment of dissipation.
Unlike grid-based methods, which typically employ Riemann solvers (Godunov 1959) to capture fluid discontinuities, SPH primarily relies on artificial viscosity (AV) formulations [^1] There have also been proposals to adapt Riemann solvers for use in SPH (Inutsuka 2002; Monaghan 1997; Cha & Whitworth 2003; Cha et al. 2010), representing a promising parallel line of development (Rosswog 2024).. The theoretical foundations of the standard AV scheme have long been established (Monaghan & Gingold 1983; Monaghan 1992). However, practical applications of the method still require additional refinements if an algorithm is desired that is sufficiently versatile to handle different fluid regimes.
One of these refinements was the introduction of corrective factors (AV limiters) that suppress viscosity in strongly shearing regions, thus reducing dissipation in vorticity-dominated flows and facilitating the growth of hydrodynamic instabilities (Balsara 1995). A widely adopted and often more effective strategy employs time-dependent dissipation coefficients (“switches”) that remain low by default, rise rapidly when a shock is detected, and relax back to their nominal values once the shock has passed (Morris & Monaghan 1997; Monaghan 2005). The different implementations of those time-dependent AV parameters primarily differ in the design and responsiveness of the shock detector and the relaxation law (Cullen & Dehnen 2010; Read & Hayfield 2012; Rosswog 2015, 2020a).
Although artificial-viscosity switches (AVSWs, hereafter) represent a major advance and do reduce unwanted dissipation, they are not without drawbacks. The formulation is largely heuristic rather than derived from first principles (though see Rosswog 2020a, for an entropy-steered variant). In post-shock regions, oscillations in the time-dependent AV coefficients can induce excessive numerical noise and spuriously increase entropy. Moreover, for certain problems, such as instability growth and subsonic turbulence, reported convergence rates are often lower than those achieved with mesh-based methods (Bauer & Springel 2012).
Recently, Frontiere et al. (2017) proposed the use of slope-limited reconstruction (SLR) within the artificial-dissipation terms to suppress spurious viscosity. Conceptually, SLR acts as a local filter that subtracts the linear component of the velocity field, preventing bulk shear from feeding the dissipative operator. In previous studies, it has been combined with conventional AV switches (Frontiere et al. 2017; Rosswog 2020a; Rosswog & Diener 2024; Cabezón et al. 2025) with good effect. Nevertheless, several tests indicate that further gains are achieved when AVSWs are omitted altogether (Sandnes et al. 2025). The prospect of attaining high accuracy without AV switches is noteworthy, and given the ubiquity of these terms in SPH, it warrants closer examination.
In this work, we carry out a suite of SPH simulations using a SPH formulation distinct from that of Sandnes et al. (2025), with two goals: (i) independently validate the performance of the slope-limited reconstruction (SLR) technique and (ii) refine it where beneficial. Our test suite includes several benchmarks considered by Sandnes et al. (2025) as well as additional problems that probe shear- and turbulence-dominated regimes. We pay particular attention to turbulent flows, where excessive artificial dissipation is especially detrimental.
This article is organized as follows. Section 2 summarizes the different AV schemes currently used to handle dissipation in, as well as our proposed methods. Section 3 presents various tests comparing the performance of these AV schemes. Section 4 is entirely dedicated to the study of subsonic turbulence, and how the AV schemes influence SPH’s ability to resolve it. The schemes that offer the best balanced results across all tests are ranked in Section 5. Finally, the Conclusions discuss the main findings of our work, the recommended AV setting, and future prospects.
2 SPH flavor
Most of the test cases described in Sect. 3 were carried out with the SPH code SPHYNX (Cabezón et al. 2017). Only the turbulence simulations were performed with SPH-EXA, owing to its scalability and native GPU acceleration. This enabled higher-resolution, time- and resource-intensive runs to be executed substantially faster. Both codes incorporate the same improved algorithms for gradient estimation (García-Senz et al. 2012; Cabezón et al. 2012; Rosswog 2015), kernel interpolation (Cabezón et al. 2008; Cabezón & García-Senz 2024), and an accurate partition of unity (García-Senz et al. 2022). These ingredients are fully consistent with the Lagrangian nature of SPH, yielding a conservative scheme. The remainder of this section focuses on the construction and implementation of dissipation in artificial viscosity–based SPH schemes.
2.1 Artificial viscosity schemes
As in other modern SPH formulations (Price et al. 2018; Rosswog 2020b; Wadsley et al. 2017), SPHYNX and SPH-EXA handle dissipation with the classical artificial viscosity (AV) scheme (Monaghan 1992, 1997). In particular, Monaghan (1997) defines the pairwise AV term $\Pi_{ab}$ in analogy with a Riemann problem,
$\Pi_{ab}=-\frac{\alpha}{2}\frac{v_{ab}^{sig}\omega_{ab}}{\bar{\rho}_{ab}}\,$, (1)
with
$\omega_{ab}={\bf v}_{ab}\cdot\hat{\bf r}_{ab}$ (2)
where ${\bf v}_{ab}={\bf v}_{a}-{\bf v}_{b}$ and $\hat{\bf r}_{ab}$ is the unit vector joining particles $a$ and $b$.
$v_{ab}^{sig}=c_{a}+c_{b}-\gamma\omega_{ab}\,$, (3)
being a characteristic signal velocity at which the information propagates between particles $a$ and $b$, and $\gamma=3$. Equations 1 and 3 are equivalent to the widely used expression of Monaghan (1992),
$\Pi_{ab}=\begin{cases}\frac{-\alpha\bar{c}_{ab}\mu_{ab}+\beta\mu_{ab}^{2}}{\bar{\rho}_{ab}}&\text{for\penalty 10000\ \penalty 10000\ }{\bf v}_{ab}\cdot{\bf r}_{ab}<0,\\ 0&\text{for\penalty 10000\ \penalty 10000\ }{\bf v}_{ab}\cdot{\bf r}_{ab}\geq 0,\end{cases}$ (4)
where
$\mu_{ab}=\frac{\bar{h}_{ab}{\bf v}_{ab}\cdot{\bf{r}_{ab}}}{{\bf r}_{ab}^{2}+\eta^{2}}\,.$ (5)
Neglecting the regularization $\eta$ and taking $\frac{\bar{h}_{ab}}{|{\bf r}_{ab}|}=1$ yields $\mu_{ab}\approx\omega_{ab}$, so comparing Eqs.1 and 4 gives $\beta=\gamma\alpha/2$. For $\gamma=3$, we have $\beta=\frac{3}{2}\alpha$. However, a more frequent choice is $\beta=2\alpha$ which better reduces post-shock oscillations in strong shocks. In the continuum limit, it has been shown that the linear part of AV in Eq. 4 is equivalent to the shear and bulk viscosities in the Navier-Stokes equations (Monaghan 2005). Thus, the AV is not as ’artificial’ as it might seem after all.
Modern AV schemes handle $\alpha$ and $\beta$ independently, since $\alpha$ could be allowed to vary per particle and in time during calculations, while $\beta$ is kept constant. For example, Wadsley et al. (2017) considers
[…]
where $\omega_{ab}$ is given by Eq. 2 and $\alpha_{min}\leq\alpha_{a}\leq\alpha_{max}$, while $\beta$ remains constant. SPHYNX and SPH-EXA also use Eq. 6 to implement AV. Typical values of the AV coefficients, when switches are used, are $\alpha_{min}=0.05$, $\alpha_{max}=1$, and $\beta=2$.
Table 1: Summary of artificial-viscosity (AV) variants considered. SLR denotes the slope-limited velocity reconstruction of Eqs. (9)–(10), optionally modulated by the Balsara factor $\mathcal{B}$ ($0\leq\mathcal{B}\leq 1$) via Eqs. (16)–(17). “Switches” indicates time-dependent $\alpha(\mathbf{r},t)$. The last row is discussed in relation to the subsonic turbulence scenario (Sec. 4). The $\beta$ AV coefficient is always fixed to $\beta=2$.
Method SLR $\mathcal{B}$ modulation Switches $\alpha$ AV coefficient Notes
1: AV No No $\alpha=1$ Classical AV baseline.
2: AVSW No Yes $\alpha_{a}\in[\alpha_{\min},\alpha_{\max}]$ Time-dependent $\alpha$.
3: AVSLR Yes No No $\alpha=1$ Slope-limited reconstruction.
4: AVSWSLR Yes No Yes $\alpha_{a}\in[\alpha_{\min},\alpha_{\max}]$ Hybrid: switches + SLR.
5: AVSLRB Yes $(1-\mathcal{B})$ No $\alpha=1$ No SLR when $\mathcal{B}\simeq 1$; strong SLR when $\mathcal{B}\simeq 0$.
[…]
2.1.1 Artificial viscosity with shock detection: switches
The choice of $\alpha$ and $\beta$, controlling the linear ($\propto\omega_{ab}$) and the quadratic ($\propto\omega_{ab}^{2}$) contributions to the AV in Eq. 6 is crucial in practical applications of SPH. The simplest option is to keep $\alpha\penalty 10000\ \mathrm{and}\penalty 10000\ \beta$ constant during the calculation. However, values of $\alpha$ large enough to capture strong shocks introduce excessive viscosity in smooth regions of the fluid. Over-dissipation damps short-wavelength modes and hinders the growth of fluid instabilities and turbulence. It also suppresses differential rotation, resulting in a poor depiction of, e.g., accretion disk dynamics.
A plausible, albeit phenomenological, remedy was proposed by Balsara (1995): adopt constant $\alpha=1$ and $\beta=2$, but multiply $\Pi_{ab}$ by a limiter ${\mathcal{B}}_{ab}$ designed to approach unity in compressive (shock) regions and to decrease toward zero in curl-dominated flows,
$\Pi_{ab}\xleftarrow{}{\mathcal{B}}_{ab}\Pi_{ab}\,;\qquad{\mathcal{B}}_{ab}=\frac{1}{2}({\mathcal{B}}_{a}+{\mathcal{B}}_{b})\,$, (7)
with
${\mathcal{B}}_{a}=\frac{|\nabla\cdot\mathbf{v}|_{a}}{|\nabla\cdot\mathbf{v}|_{a}+|\nabla\times\mathbf{v}|_{a}+10^{-4}\frac{c_{a}}{h_{a}}}\,$, (8)
where $c_{a}$ and $h_{a}$ are the speed of sound and the smoothing length of particle $a$, respectively.
Later, a more efficient algorithm was introduced by Morris & Monaghan (1997) and subsequently refined by others. In this approach, the linear AV coefficient is kept near a low default value, $\alpha=\alpha_{min}$ (with $\alpha_{min}\simeq 0.05$), but is quickly increased toward $\alpha_{max}$ when a shock is detected. Cullen & Dehnen (2010) showed that a fast and robust shock detection is crucial in this technique, and proposed using a higher-order indicator based on the derivative of $\nabla\cdot{\bf v}_{a}$, which generally yields improved results. Several variations and refinements of these AV switches for controlling $\alpha({\bf r},t)$ can be found in the literature (Read & Hayfield 2012; Rosswog 2015).
At this point, it is worth highlighting the subtle difference between a limiter and a switch in this context. A limiter, like Balsara’s, is a continuous, instantaneous, memoryless amplitude control that keeps AV high by default, except in regions that meet well-defined physical criteria (e.g. a high curl value). A switch, on the other hand, is a tunable sensor with temporary memory that can lag or overshoot in different regions. Switches are engineered to work well in some tests, but their exact form is a practical choice rather than a unique consequence of the equations, hence their heuristic.
Despite their success, AVSWs can introduce additional noise, especially in the post-shock wake where $\alpha$ may oscillate due to rapid changes in the shock indicator, leading to spurious entropy production (e.g. Cullen & Dehnen 2010; Read & Hayfield 2012; Rosswog 2015). In mixed regimes where shocks and strong shear coexist, for example when a dilute stream impacts a dense cloud, switches can struggle to balance compression-driven dissipation against shear suppression (cf. Balsara 1995; Price et al. 2018). Finally, switches tend to be less effective at controlling low-level numerical noise in otherwise quiescent regions; dedicated “noise triggers” can be added to mitigate this, but at the expense of additional heuristics and computing time (Rosswog 2015).
Table 2: Simulations name convention and some features of the studied test cases. The first column is the test type; the second to seventh columns show the test name that relates the test type to the methods defined in Table 1. The three last columns indicate the size of the computational box, the total number of SPH particles, and the equivalent number per axis, N1D in a cubic box with $L=1$, calculated as $N^{\mathrm{1D}}=\left[\frac{N}{L_{x}\cdot L_{y}\cdot L_{z}}\right]^{1/3}$. The second AV coefficient is chosen $\beta=2$ in all tests
Test AV AVSW AVSLR AVSWSLR AVSLRB $AVSLRB2{\mathcal{F}}$ Box NN $N^{\mathrm{1D}}$
$\alpha=1$ $\alpha(t)$ $\alpha=1$ $\alpha(t)$ $\alpha=1$ $\alpha=1$ $\alpha=1$, ${\mathcal{F}}=0.5$ $L_{x},L_{y},L_{z}]$ $(\times 10^{6})$
Shock-Tube ST1 ST2 ST3 ST4 ST5 ST6 ST7 $[1,1/8,1/8]$ 0.97 396
Sedov S1 S2 S3 S4 S5 S6 S7 $[1,1,1]$ 8.0 200
GH-Vortex V1 V2 V3 V4 V5 V6 V7 $[1,1,0.5]$ 0.5 100
KHI KH1 KH2 KH3 KH4 KH5 KH6 KH7 $[1,1,0.1]$ 0.6 182
RTI RT1 RT2 RT3 RT4 RT5 RT6 RT7 $[1,2,0.1]$ 1.2 182
Wind-Cloud WC1 WC2 WC3 WC4 WC5 WC6 WC7 $[40,10,10]$ 0.5 200
Turbulence T1 T2 T3 T4 T5 T6 T7 $[1,1,1]$ 512 800
Figure 1: Velocity profile, $v_{x}$, for the shock-tube test and the models shown in the second row in Table 2, at time $t=0.2$
2.1.2 An artificial viscosity algorithm free of switches
A linear velocity field implies a smooth flow without the presence of shocks. Therefore, linear velocity fields should have a vanishing AV. Following the prescription introduced by Frontiere et al. (2017), we reconstruct the local velocity field for each particle by removing the linear component of the velocity jump that is used to trigger and calculate the AV (Eq. 6). Assuming differentiability in the neighborhood of each SPH particle, the optimal linear approximation of the velocity field is given by its Jacobian acting on the separation vector. Rather than using $\mathbf{v}_{ab}=\mathbf{v}_{a}-\mathbf{v}_{b}$ in the AV, we use $\mathbf{v}^{\prime}_{ab}=\mathbf{v}^{\prime}_{a}-\mathbf{v}^{\prime}_{b}$, defined as:
$v^{\prime}_{a,i}$ $\equiv v_{a,i}-\frac{1}{2}\phi_{ab}\mathbf{J}_{\mathbf{v}_{a}}\mathbf{x}^{T}_{ab}\,$, (9)
$v^{\prime}_{b,i}$ $\equiv v_{b,i}+\frac{1}{2}\phi_{ba}\mathbf{J}_{\mathbf{v}_{b}}\mathbf{x}^{T}_{ab}\,$, (10)
where $\mathbf{x}^{T}_{ab}=\mathbf{x}_{a}-\mathbf{x}_{b}$ is the column vector of the relative distance between particles $a$ and $b$, and $\mathbf{J}_{\mathbf{v}}$ is the Jacobian of the velocity field. In the presence of discontinuities, such a linear estimator is not accurate enough; hence, the factor $\phi_{ab}$ is designed to approach zero in these cases and to remain close to unity for smooth flow. Frontiere et al. (2017) proposed a van Leer-like limiter (van Leer 1974):
$\phi_{ab}$ $=\max\left[0,\min\left(1,\frac{4F_{ab}}{(1+F_{ab})^{2}}\right)\right]\times\kappa_{ab}\,$, (11)
[…]
$q_{ab}$ $\equiv\min(q_{a},q_{b})\,$, (13)
$q_{crit}$ $\equiv\left(\frac{32\pi}{3nb_{a}}\right)^{1/3}\,$, (14)
$F_{ab}$ $\equiv\frac{\mathbf{x}_{ab}\mathbf{J}_{\mathbf{v}_{a}}\mathbf{x}^{T}_{ab}}{\mathbf{x}_{ab}\mathbf{J}_{\mathbf{v}_{b}}\mathbf{x}^{T}_{ab}}\,$, (15)
where $q_{a}=|\mathbf{x}_{b}-\mathbf{x}_{a}|/h_{a}$, $q_{crit}$ estimates the mean interparticle spacing in units of smoothing length, and $nb_{a}$ is the neighbor count of particle $a$. The parameter $q_{fold}$ sets the width of the Gaussian correction (we adopt the recommended value from Frontiere et al. (2017) of $q_{fold}=0.2$).
Equations 9 and 10 remove the linear component of the velocity field, limiting the slope at discontinuities and reducing unwanted dissipation in smooth flows. We refer to this slope-limited reconstruction as AVSLR, and contrast it with traditional AVSW (time-dependent switch) methods in what follows.
Currently, few calculations incorporating these AVSLRs have been reported (Frontiere et al. 2017; Rosswog 2020a, b; Cabezón et al. 2025; Sandnes et al. 2025), and they were mostly used in combination with switches. As in Sandnes et al. (2025), our premise is that AV switches are unnecessary when using the slope-limited reconstruction algorithm. We therefore explore several variants of the implementation of the SLR (see Table 1):
  • AVSLR: Use Eqs. 9, 10 to remove the linear component of the velocity field.
  • AVSLRB: modulate the SLR with the Balsara limiter:
    $v^{\prime}_{a,i}$ $\equiv v_{a,i}-\frac{1}{2}\left(1-{\mathcal{B}}^{p}_{a}\right)\phi_{ab}\mathbf{J}_{\mathbf{v}_{a}}\mathbf{x}^{T}_{ab}\,$, (16)
    $v^{\prime}_{b,i}$ $\equiv v_{b,i}+\frac{1}{2}\left(1-{\mathcal{B}}^{p}_{b}\right)\phi_{ba}\mathbf{J}_{\mathbf{v}_{b}}\mathbf{x}^{T}_{ab}\,$, (17)
    with $\mathcal{B}_{a,b}$} defined by Eq. 8. Following the suggestion of Sandnes et al. (2025) we first consider the case with $p=1$.
  • AVSLRB2 same as above but with $p=2$; this does not change the description of shocks, although it proved more advantageous than $p=1$ in our tests.
  • Baselines: for comparison, we also consider (i) classical AV with fixed coefficients (AV), (ii) AV with time-dependent switches (AVSW), and (iii) a hybrid combining switches with slope-limiting (AVSWSLR).
The objective is to find the most balanced and general scheme among all these options in light of the results.
3 TESTS
Figure 2: Evolution of maximum density in the Sedov test for the different models described in Table 2.
We first consider six test problems designed to compare the performance of the proposed schemes across shock- and shear-dominated regimes. The shock-dominated set comprises the Sod shock tube (Sod 1978) and the Sedov blast (Sedov 1959). The shear/instability set comprises the steady Gresho–Chan vortex (Gresho & Chan 1990), Kelvin–Helmholtz (KH) and Rayleigh–Taylor (RT) instabilities, and the wind–cloud interaction in which a dilute supersonic wind disrupts a dense cloud (Agertz et al. 2007). A seventh test dealing with subsonic turbulence is deferred to Sect. 4. When needed, we adopt an ideal-gas equation of state $P=(\gamma-1)\rho u$, with $\gamma=5/3$ (and $\gamma=1.4$ for the RT test). Simulations were carried out with SPHYNX [^2] https://astro.physik.unibas.ch/sphynx (Cabezón et al. 2017), including the updates discussed in García-Senz et al. (2022). Kernel interpolation uses the mixed sinc kernel $S^{0.9}_{4,9}$ (Cabezón & García-Senz 2024), which is pairing-resistant and robust for a wide range of neighbors, $60\leq n_{b}\leq 400$. Unless stated otherwise, we set $n_{b}=100$. Additional setup details for each case are summarized in Tables 1 and 2.
To make quantitative comparisons between models, we report a mean absolute error $L_{1}$ wherever a suitable reference is available:
$L_{1}=\frac{1}{N_{\mathcal{S}}}\sum_{i\in\mathcal{S}}\bigl|p^{\rm SPH}_{i}-p^{\rm ref}_{i}\bigr|\,$, (18)
where $S$ denotes the set of sampling points with positions $d_{i}$ in the range $d_{0}\leq d_{i}\leq d_{f}$ (e.g., a 1D line-out or radial profile), and $N_{S}$ is the number of particles belonging to the set $S$. Here $p^{\rm SPH}$ is the measured quantity from the simulation and $p^{\rm ref}$ is the corresponding reference (analytic solution or high-resolution benchmark), specified for each test below.
3.1 The shock-tube
In this test, a box with the size and number of equal-mass particles specified in Table 2 is initialized with a gas that has a strong discontinuity in density and pressure across a contact interface at $(x=x_{c})$. The problem has a well-known analytical solution (Sod 1978), which we use to assess the performance of the different artificial-viscosity (AV) prescriptions.
Figure 3: Density and radial velocity profiles in the Sedov-Taylor test at time $t=1$ for the different models described in Table 2. The solid line represents the analytical reference. The $L_{1}$ error between the numerical and analytical estimates from the center until the maximum radial velocity is also shown.
The setup follows common practice (Wadsley et al. 2017; Price et al. 2018): on the left state $(x<x_{c})$, we take $\rho_{l}=1$, $p_{l}=1$; on the right state $(x>x_{c})$, $\rho_{r}=0.125$, $p_{r}=0.10$. The particle distribution across the contact was left unsmoothed, but the pressure was smoothed according to
$P_{0}=-A\penalty 10000\ \tanh\left(\frac{x-x_{c}}{\delta}\right)+B\,$, (19)
with $\delta=0.01$, $x_{c}=0.5$, $A=0.45$, and $B=0.55$.
This test probes how the different AV choices handle strong shocks. The results are summarized in Fig. 1, which shows the $x$-velocity profile, $v_{x}$. We compute the $L_{1}$ error using Eq. 18 over $x\in[x_{0},x_{f}]$ with $x_{0}=d_{0}=0.05$ and $x_{f}=d_{f}=0.868$. The upper limit $x_{f}=0.868$ is chosen to exclude the terminal vertical drop in the velocity so that the sum focuses on the resolved shock, contact, and rarefaction features.
From Fig. 1, models ST2 and ST4, those employing switches, perform worst, exhibiting pronounced velocity oscillations in the post-shock region. Substantially better results are obtained either with the simple, constant-parameter AV (ST1), or with the SLR approach, modulated with the Balsara limiter (ST5 and ST6). The ST3 calculation is intermediate, with oscillations confined to the immediate shock vicinity. Overall, the best performance in this set is delivered by ST1, followed by ST6, which uses SLR and applies the squared ($p=2$) Balsara limiter to the functions in Eqs. 16 and 17.
3.2 Spherical explosion
Figure 4: Azimuthal velocity distribution in the Gresho-Chan vortex experiment at $t=1$ for models Vk in the third row in Table 2. The analytical reference is indicated by the continuum line.
Usually referred to as the Sedov-Taylor blastwave, we perform this test in a three-dimensional box of side $L=1$ with $N=200^{3}$ particles arranged in a glass-like configuration. All particles have the same mass, resulting in a homogeneous initial density profile $\rho=1\pm 0.005$. The explosion is initiated at the box center by depositing a total energy $\Delta U=1$ into the gas, distributed with a Gaussian profile of characteristic width $\delta=2h$, where $h$ is the smoothing length at $t=0$.
Figures 2 and 3 summarize the results. The time evolution of the maximum density is shown in Fig. 2. Models S3 and S4 overshoot the strong-shock compression limit $\rho/\rho_{0}=(\gamma+1)/(\gamma-1)=4$ for $\gamma=5/3$, while the remaining four models stay below and approach this limit from below as the blast develops. As in the shock-tube test, we compute the $L_{1}$ error. In this case, for the radial velocity between the origin and the r coordinate of the maximum velocity. The lower panels of Fig. 3 show the profile of $v_{rad}$ and the $L_{1}$ value at $t=0.1$. In the central region ($r\leq 0.25$), the numerical profiles depart from the analytical solution (solid line), though the amount of mass involved is very small. To adequately resolve the diluted central region, a much larger number of SPH particles would be needed. In the post-shock region, all models show velocity oscillations, which are particularly pronounced near the shock front for models S3 and S4, consistent with the behavior observed in the shock-tube test.
According to these results, the best models are S6, S1, and S5 (these three have very similar results), followed by S2. The worst results arise when using switches together with the SLR technique (S4), or with SLR alone without Balsara’s modulation (S3).
3.3 The Gresho-Chan vortex
Simulation of a pure fluid vortex (Gresho & Chan 1990) is challenging for SPH schemes working with standard AV: excessive viscosity damps the differential rotation and degrades the velocity field. The flow is axisymmetric and defined by
[…]
with $\psi=5r$, $r=\sqrt{x^{2}+y^{2}}$, $v_{0}=1$, $\rho_{0}=1$, and $P_{0}=5$.
The box size and number of SPH particles are listed in Table2. Because the initial azimuthal velocity is piecewise linear, this test is expected to favor AVSLRs methods over AVSWs. Results are summarized in Fig. 4, which also provides the $L_{1}$ error for quantitative comparisons. We compute $L_{1}$ using Eq. 18 over $r\in[d_{0},d_{f}]=[0,0.5]$, with $p^{SPH},p^{ref}$ taken as the azimuthal velocity from the SPH calculation and the analytic $v_{\varphi}$ in Eq. 20, respectively.
Figure 4 clearly shows the superior performance of AVSLRs schemes over AVSWs. As expected, the worst case is V1, which uses traditional AV with constant coefficients $\alpha=1$, $\beta=2$, and no limiter. Introducing switches yields a substantial improvement, but still falls short of the SLR methods. For this test, the best results are obtained with V4 (AVSRL), V5 (AVSLRB), and V6 (AVSLRB2), with practically no differences between them.
Figure 5: Color-coded density for the KH models in Table 2 at $t=3$ ($t/t_{\mathrm{KH}}=2.8$).
Figure 6: Amplitude mode growth of the KH models. A zoom-in of the nearby lines KH3, KH4, KH5, and KH6 around $t=1.5$ is shown in the subfigure.
3.4 The Kelvin-Helmholtz instability
The evolution of a shear layer between two fluids of different densities naturally triggers the Kelvin-Helmholtz instability (KHI). Because this instability is ubiquitous in nature, including astrophysical phenomena, its accurate simulation is of great importance. Since AV can artificially dampen KHI growth, this problem is a sensitive test for the SLR algorithm.
We initialize SPH particles in a thin three-dimensional box of the size listed in Table 2, with a 2:1 particle number ratio between the high- and low-density regions. The initial setting consists of three stratified layers, with the central layer having the higher density.
Figure 7: Colormaps of density for the RT models in Table 2 at $t=6$.
The initial number of neighbors is $n_{b}=120$. The initial velocity distribution follows McNally et al. (2012), where shear was induced by imposing $v_{x}=\pm 0.5$ across the separation layer, exponentially damped along the $Y$-direction over a characteristic length $L=0.025$. The density across the contact layer is not smoothed. We seed a single-mode normal perturbation in the cross-stream velocity,
$v_{y}(t=0)=\Delta v_{y}\sin\left(\frac{2\pi}{\lambda}x\right)\,$, (22)
where $\lambda=0.5$ and we adopt an initial background pressure of $P_{0}=2.5$.
Times are also reported normalized to the characteristic KHI growth time $t_{\mathrm{KH}}$, as defined in Agertz et al. (2007), resulting in $t_{\mathrm{KH}}=1.06$ for our models. The results are shown in Figs. 5 and 6, and Table 3. Figure 5 shows the density colormaps of the different models at time $t=3$ ($t/t_{\mathrm{KH}}=2.8$). At that time, the instability is fully nonlinear, with well-developed swirls, billows, and secondary structures. These extend over a wide range of wavelengths, with a cut-off set by the real viscosity but, in practice, by resolution and numerical dissipation.
Table 3: KH: Perturbation amplitude $A(t)$ at $t=1.5$ for the models in Table 2. The reference value from McNally et al. (2012) is $A(t=1.5)=14.79\times 10^{-2}$
- KH1 KH2 KH3 KH4 KH5 KH6
[…]
According to Fig. 5, models KH3 to KH6 develop secondary structure more robustly than KH1 (traditional AV with constant coefficients and no switches). Steering dissipation with switches (KH2) is the second worst. This is quantified in Fig. 6, which displays the evolution of the perturbation amplitude. KH3 and KH6 show the fastest growth. At $t=1.5$, the amplitudes (presented in Table 3) are close to the reference (McNally et al. 2012), which is remarkable given the moderate resolution in these simulations $N^{1D}=182$.
We also see that the use of switches (KH2), represents a considerable improvement over traditional AV with constant coefficients (KH1). Even so, the computed growth rate is much lower than that obtained with the SLR family, especially KH3 and KH6. The KH5 and KH4 models show slightly less amplitude growth than KH3, but still larger than KH2, and can be considered good options when balancing accuracy and robustness.
3.5 The Rayleigh-Taylor instability
The study of convective-like instabilities in a gravitational field is of major importance in astrophysics. Historically, SPH has struggled to model these systems. In particular, the growth of the Rayleigh-Taylor instability (RTI) was often poorly depicted because AV introduces excessive damping and kills the smallest perturbation modes. However, combining improved gradient estimators with AV switches made a reasonable approach to this phenomenon possible (Cabezón et al. 2017; Rosswog 2020a).
Table 4: RTI: Kinetic energy (KE) in the Y-direction at $t=6$ for the models in Table 2.
- RT1 RT2 RT3 RT4 RT5 RT6
KE ($\times 10^{-3}$) 6.81 8.49 8.87 8.39 8.62 8.87
In this test, we use a box of size $[1,2,0.1]$ filled with a two-density fluid. The upper half has $\rho_{u}=2$, whereas the lower half $\rho_{d}=1$. The system is immersed in a constant gravitational field with g=0.5g=-0.5, and the initial pressure profile is set so that the hydrostatic balance is satisfied. The interface between the low- and high-density regions is sharp (no smoothing). The pressure at the bottom boundary is $P(y=0)=2.5$. The pressure profile is
[…]
where $y_{b}$ is the interface location, $\bar{\rho}$ is the average density around $y_{b}$, and $\delta=0.4$ is the initial width of the unstable layer $y_{b}-\delta/2$, $y_{b}+\delta/2$]. The fluid is initially at rest, and we seed a velocity perturbation $\Delta v_{y}$ localized to the interface,
$v_{y}=\Delta v_{y}\left[1-\cos\left(2\pi x\right)\right]\left[1+\cos\left(\frac{2\pi(y-y_{b})}{\delta}\right)\right]\,,\ \mathrm{for}\ |y-y_{b}|\leq\frac{\delta}{2}\,$, (24)
with $\Delta v_{y}=0.025$.
Despite the moderate resolution $N^{\mathrm{1D}}=182$, we can assess performance by inspecting the density maps in the non-linear regime and the kinetic energy in the vertical direction. Figure 7 shows the color-coded density at $t=6$, while Table 4 provides quantitative information on the kinetic energy stored in vertical displacements at the same time. The RTI grows and becomes nonlinear even when with traditional AV using constant $\alpha=1$, $\beta=2$ (RT1), but without discernible substructure. Introducing a time-dependent $\alpha$ (RT2) significantly improves the outcome, allowing the development of secondary structures and larger bubble/spike amplitudes. Activating the SLR (RT3) produces a similar qualitative improvement. However, although the AVSW calculation develops structure comparable to the AVSLR models, the stored kinetic energy in the vertical direction is clearly lower (Table 4). Among the SLR variants, RT3, RT5, and RT6 are similar, with RT3 and RT6 exhibiting the highest kinetic energy, and thus preferred, closely followed by RT5.
Model RT4, which combines switches and SLR, reduces AV via both mechanisms. The resulting lower dissipation amplifies numerical noise, leading to more asymmetric and more diffuse substructures. Despite the reduced viscosity, its vertical kinetic energy is the second lowest in the set, reflecting the detrimental impact of noise on coherent RT growth.
3.6 The Wind-Cloud test
The destruction of a dense cloud embedded in a dilute supersonic wind is demanding test that requires an accurate treatment of both shocks and shear-driven instabilities (Agertz et al. 2007, and references therein). That work showed that traditional SPH underperformed relative to mesh methods, largely due to tensile instability that permeated SPH codes at the time. Modern SPH schemes, however, can handle this scenario well and yield competitive results (Frontiere et al. 2017; Wadsley et al. 2017; García-Senz et al. 2022). While for this problem the mitigation of E0-errors is more decisive than the choice of AV, the AV prescription still modulates the outcome and merits discussion.
The initial configuration follows Frontiere et al. (2017). The box size and particle count are given in Table 2. The wind and cloud have initial densities $\rho=1$ and 10, respectively, and are in pressure equilibrium at $P_{0}=1$. The cloud is initially at rest, while the wind moves supersonically with Mach number $\mathcal{M}=2.7$. Wind and cloud particles are distributed in glass-like lattices, and the interface is sharp.
Numerical studies have identified different hydrodynamic regimes as a function of the KHI characteristic time $t_{KH}=2.9$ (coresponding to our setup [^3] We take the characteristic $t_{KH}$ time as in Agertz et al. (2007), which is defined as $t_{KH}=1.6\penalty 10000\ t_{crush}$, where $t_{crush}$ is the cloud crushing time, right after the formation of the bow-shock; beyond this, KHI erodes the cloud surface.) at which the KHI dominates:
  • $t\lesssim t_{KH}$. Direct impact and bow-shock formation control the initial stripping of the material from the sphere. According to Fig. 8, the cloud evolution is largely insensitive to the AV scheme.
  • $t_{KH}\lesssim t\lesssim 3t_{KH}$. Vorticity, primarily from KHI, peels and mixes cloud material into the wind. As seen in Fig 8, the removal occurs at a remarkably steady rate. The destruction rate $\mathcal{R}=\dot{M}(t)/M_{0}$ differs across AV schemes and provides a convenient ranking criterion (Table 5).
  • $t\geq 3t_{KH}$ Complete dissolution of the cloud is challenging to model physically and is also resolution dependent, since mass remnants approach zero at sufficiently high resolution. In SPH, a small artificial thermal conductivity is commonly added to achieve complete mixing and thermodynamic equilibrium (Price 2008). By default, SPHYNX includes an artificial heat conduction term, following Price et al. (2018) with a diffusion coefficient $\alpha_{u}=0.1$.
Overall, all AV schemes aim for complete cloud destruction at $t\leq 4t_{KH}$, in reasonable agreement with the results obtained with state-of-art mesh-codes such as ENZO (cited in Agertz et al. 2007). This reflects the strengths of our SPH flavor, namely, accurate gradient estimation via the integral approach (García-Senz et al. 2012) and the geometric density average force estimation (Wadsley et al. 2017; García-Senz et al. 2022). Still, the particular implementation of the AV is also important as it modulates the path to complete dissolution. The interval $1.25\penalty 10000\ t_{KH}\lesssim t\lesssim 2.5\penalty 10000\ t_{KH}$ is particularly relevant to our study because the slope of the $M(t)/M_{0}$ curves allows an easy classification of the AV variants.
According to Table 5, the highest destruction rates at $1.25\penalty 10000\ t_{KH}\leq t\leq 2.5\penalty 10000\ t_{KH}$ occur for WC6, while the slowest are those with bare AV or with switches alone. Using this criterion, the AV schemes are ranked, from worst to best, in the same order as they appear in the table (we use the first row for ranking). However, note that SLR models without Balsara modulation (models WC3 and WC4) take longer to dissolve the cloud because they overestimate the effect of shocks, as seen in the Sedov test (Fig. 2), which in the wind-cloud problem delays efficient KHI-driven shredding.
Table 5: Cloud destruction rate, $\mathcal{R}=\dot{M}/M_{0}$, at $t=1.8\ t_{KH}$ and $t=2.0\ t_{KH}$ for the WC models.
WC1 WC2 WC3 WC4 WC5 WC6
$\mathcal{R}_{1.8t_{KH}}$ -0.105 -0.112 -0.139 -0.144 -0.152 -0.162
$\mathcal{R}_{2.0t_{KH}}$ -0.121 -0.119 -0.127 -0.152 -0.174 -0.214
Figure 8: Evolution of the cloud mass normalized to its initial value. The elapsed time is normalized to $t_{KH}=2.9$.
4 Subsonic turbulence
Turbulence is ubiquitous across physical scales and is especially relevant in astrophysics, where it plays a major role in chemical mixing, stellar collapse, and supernova explosions among others. High-resolution simulations of turbulence are delicate: they are long and computationally expensive, and accuracy must be maintained over many timesteps. Subsonic turbulence is particularly demanding: even small accumulations of numerical error or a modest excess of dissipation can suppress the development of a faithful inertial range (i.e., the Kolmogorov cascade down to the dissipation scale). This has long challenged SPH, which historically lagged behind Eulerian methods at low Mach numbers, until recently.
Modern implementations that combine accurate gradient estimators, generalized volume elements, pairing-resistant kernels, and tighter AV control have now achieved Eulerian-quality results (Cabezón et al. 2025). The simulation of subsonic turbulence presents an inverted dissipation problem relative to the previous tests in this work. In shocks, disks, and hydrodynamical instabilities, dissipation is driven primarily by strong compression and large velocity gradients. In contrast, subsonic turbulence is limited by the background numerical dissipation floor: if this baseline dissipation is even slightly too high, the energy cascade is prematurely damped, and the inertial range collapses. To mitigate spurious triggering by small-scale particle noise, it is therefore necessary to control the amplitude of the dissipation. To this end, we reuse the Balsara coefficients (already employed to modulate the SLR terms) to multiply the linear part of the AV, clamped from below by a floor $\mathcal{F}$:
$\Pi_{ab}=\frac{-\max({\mathcal{F}},\bar{\mathcal{B}}_{ab})\alpha\bar{c}_{ab}\mu_{ab}+\beta\mu_{ab}^{2}}{\bar{\rho}_{ab}}\,,\ \text{for\penalty 10000\ \penalty 10000\ }{\bf v}_{ab}\cdot{\bf r}_{ab}<0\,$, (25)
with $0\leq{\mathcal{F}}\leq 1$ and $\bar{\mathcal{B}}_{ab}=\frac{1}{2}({\mathcal{B}}_{a}+{\mathcal{B}}_{b})$. The clamp reflects our goal of a general-purpose dissipation scheme: a floor value $\mathcal{F}=1$ recovers the scheme used in previous tests. Allowing ${\mathcal{F}}\rightarrow 0$ yields excellent results in the subsonic turbulence, but at the cost of effectively disabling AV in shear-dominated regions adjacent to shocks or in mixed regimes. We adopt the compromise of $\mathcal{F}=0.5$: AV is sufficiently attenuated for subsonic turbulence, yet does not vanish in, e.g., shock-dominated flows. According to Table 7, this approach produces good results in all tests considered in this work.
Our metric to rank the turbulence models is to measure the location of the bottleneck peak ($k_{max}$) in the compensated velocity power spectrum. The works of Valdarnini (2016) and Cabezón et al. (2025) showed that SPH can reproduce both the inertial range and the bottleneck effect, provided that accurate gradients and sensitive AV control are used jointly. Therefore, given such a modern SPH implementation and a determined resolution, $k_{max}$ becomes a proxy for effective dissipation: without amplitude control of the AV, dissipation acts earlier (lower $k$) and the bottleneck intrudes sooner, effectively reducing the inertial range. Hence, a larger $k_{max}$ indicates a potential better model in this context.
Given the computational cost, we used the GPU-native SPH-EXA code [^4] https://github.com/sphexa-org/sphexa. to run at relatively high resolution, using $N=800^{3}$ particles. The domain is a box of side $L=1$ with periodic boundary conditions and constant density $\rho=1$. Particles start from a relaxed glass-like distribution. The driving is done via an Ornstein-Uhlenbeck process (Uhlenbeck & Ornstein 1930; Federrath et al. 2010), following a parabolic amplitude field $\sim 1-(k-k_{c})^{2}$, with $k=|\mathbf{k}|$, and $k_{c}=2$, which corresponds to a length $L/2$, and vanishing for $k\leq 1$ and $k\geq 3$. This approach is implemented to reduce the impact of the stirring on the energy cascade, facilitating a natural transfer of energy from larger to smaller scales. Our forcing uses a natural mixing of $1/3$ compressive and $2/3$ solenoidal contributions. All simulations begin in a homogenous state and are stirred until they become statistically stationary at $t\simeq 10$ with an RMS Mach number of $\mathcal{M}=0.3$. The times are normalized to the sound-crossing time, which in our setup is $t_{sc}=L/c=1$. We used a quasi-isothermal EOS with adiabatic index $\gamma=1.001$. This setup is provided out-of-the-box in SPH-EXA and can be reproduced by modifying the AV as desired and using the following command line:
\$ ./sphexa-cuda --init turbulence --prop turbulence -n 800 -s 10.0 -w 1.0
where -n sets $N^{1D}$, -s is the simulated physical time, and -w is the output frequency.
Figure 9: Instantaneous compensated velocity power spectra at $t=10$.
Table 6: Subsonic turbulence test: $k$ values for the maximum of the power spectra of the velocity field at the bottleneck.
T1 T2 T3 T4 T5 T6 T7
$k_{max}$ 8.68 25.40 46.88 59.04 32.99 41.67 49.47
Figure 9 shows the resulting compensated power spectra at $t=10$, and Table 6 lists the corresponding $k_{max}$ values for models $T_{k}$ in Table 2. As expected, AV with constant coefficients (T1) activates too early, pushing the bottleneck effect to low $k$, near the box scale, leaving no inertial plateau and quickly leading to viscous roll-off. The best model combines AVSW with SLR (T4) yielding nearly a decade-wide plateau and bottleneck peaking at $k=59$. Pure SLR-based models, without AV amplitude modulation, achieve a very similar inertial range as (T4), meaning that the large-to-intermediate dynamics are comparable. However, they show an earlier (and slightly stronger) encroachment of the bottleneck into the dissipation range. This is not surprising, since there is no mechanism to suppress AV in a smooth flow. Any residual compression/noise from particle disorder keeps a finite $\alpha$ active essentially all the time, leading to the earlier onset of dissipation. Model T7 addresses this point, limiting the linear component of the AV, which delivers the second-best results.
5 Searching for the optimal AV scheme
Table 7: Summary of the quantitative information used in Fig. 10 to rank the first six tests cases with six AV schemes in Table 2. The same information is also shown in the last row for a configuration giving good results in the turbulence study (see Sect. 4)
Test ST S V KHI RTI WC T
$\mathrm{L}_{1}(v_{x})$ $\rho_{max}$ $\mathrm{L}_{1}\ (v_{rad})$ $\mathrm{L}_{1}(v_{\phi})$ $\mathrm{A}(t/t_{KH}=1.5)$ $\mathrm{KE_{Y}}$ $\mathcal{R}_{1.8t_{KH}}$ $k_{max}$
$(\times 10^{-3})$ $(\times 10^{-2})$ $(\times 10^{-2})$ $(\times 10^{-2})$ $(\times 10^{-3})$
1-AV 2.97 4.00 2.57 5.80 2.87 6.81 -0.105 8.68
2-AVSW 5.90 4.02 2.69 2.54 8.31 8.49 -0.112 25.40
3-AVSLR 3.16 4.54 3.34 1.22 12.58 8.87 -0.139 46.88
4-AVSWSLR 6.24 4.56 3.33 2.50 12.14 8.39 -0.144 59.04
5-AVSLRB 3.05 4.02 2.68 1.22 12.01 8.62 -0.152 32.99
6-AVSLRB2 3.04 4.03 2.54 1.22 12.52 8.87 -0.162 41.67
[…]
Figure 10: Ranking heatmap of all AV methods for each test. Numbers show the relative difference to the best metric result overall for each test, using the values reported in Table 7. Lower is better.
Given the spread of the results across the AV schemes in Table 2, the best scheme is the one that is most balanced over the entire test suite. Generally speaking, for shocks, the best schemes maintain a high AV at the shock front (e.g. models S1 and ST1 in Table 2). The opposite is true in regions subjected to strong shear, where AV should be kept low. On another note, the limitation in subsonic turbulence comes from the background random particle noise, which can spuriously trigger unnecessary AV. This needs to be controlled to allow for the emergence of the smaller-scale structure and the delay of the bottleneck and dissipation regimes.
Figure 10 shows a ranking heatmap of AV variants across all tests, including turbulence. The heatmap is indicative and consistent with the case-by-case discussion above and summarized in Table 7. The x-axis lists the tests, while the y-axis lists the AV schemes, indexed as in Table 7. Each cell is colored on a 0-100% scale representing the relative distance to the best metric result on that test. Quantitative metrics are those reported in the previous sections: L1 values (ST, S, V), the perturbation growth rate (KHI), the vertical kinetic energy (RTI), the cloud-destruction rate in a fixed time interval (WC), and the bottleneck peak wavenumber in the compensated velocity power spectra (T). In each column (test), the best results have a value of 0%.
[…]
6 Conclusions
Recently, Sandnes et al. (2025) suggested that AV switches may be unnecessary if a slope-limited reconstruction (SLR) is used (Frontiere et al. 2017; Rosswog 2020a, b, see also). If confirmed in general, this would represent a paradigm shift in the way AV is implemented in SPH codes.
There is, however, no free lunch. High, constant AV is excellent for shock capturing, but overdamps turbulence, suppresses the emergence of hydrodynamic instabilities, and erases small-scale shear. Low, constant AV is great for turbulence but produces strong post-shock oscillations. Switches may lag or trigger prematurely and leave residual noise in quiescent regions. Traditional Balsara limiters help in vortical flows, but they are unable to totally suppress unwanted dissipation in rotating shear flows. SLR reduces false compression triggers, but it does not provide the adequate dissipation in strong shocks and fails to describe them. Furthermore, SLR alone it is not enough to totally control spurious AV triggers by random particle movements in smooth flows. There is always a trade-off.
Even so, SLR is promising, and in our tests, it shows better results overall. In this work, we explore several SLR-based implementations across a broad test suite (shocks, shear flows, canonical instabilities, and subsonic turbulence) and suggest a balanced AV algorithm that performs robustly in all regimes.
Our main findings are:
  • AVSLR schemes outperform traditional AVSW across all tests. This is especially relevant for the study of subsonic turbulence, where SPH has traditionally performed worse than mesh-based methods.
  • In line with Sandnes et al. (2025), a good multi-regime performance requires modulating the SLR reconstruction with, for example, the Balsara limiter $\mathcal{B}$, as shown in Eqs. (16) and (17). While Sandnes et al. (2025) used the exponent $p=1$, we find that $p=2$ (our AVSLRB2) consistently yields better results, particularly in turbulence, without sacrificing shock handling.
  • Combining switches with SLR (AVSWSLR) tends to degrade shock and Gresho–Chan performance, but it works very well for subsonic turbulence, Fig. 9 and Cabezón et al. (2025). This improvement is due to the additional reduction of the AV parameter $\alpha$ introduced by the switches.
  • Motivated by the turbulence gains from reduced $\alpha$, we tested a simple amplitude control mechanism of AV in our AVSLRB2 scheme, considering $\alpha\rightarrow\alpha\times max[\mathcal{F},\mathcal{B}]$, with $0\leq{\mathcal{F}}\leq 1$ in Eq. 25. Taking $\mathcal{F}=1$ recovers the AVSLRB2 scheme, while ${\mathcal{F}}=0.5$ improves the turbulence spectra while producing only minor changes in the other tests (see the last two rows of Table 7 and Fig. 10). This offers a switch-free path to near-optimal turbulence behavior without compromising shocks or instabilities.
In short, AVSLRB2 (SLR with Balsara modulation at $p=2$) provides the best overall balance across shock-, shear-, and instability-dominated regimes, and adding a modest AV amplitude floor (${\mathcal{F}}=0.5$) recovers much of the turbulence advantage seen with switches+SLR, without the broader downsides of switch-driven dissipation control.
[…]
Acknowledgements.
This work has been supported by the Swiss Platform for Advanced Scientific Computing (PASC) project ”SPH-EXA: Optimizing Smoothed Particle Hydrodynamics for Exascale Computing”. It has also been carried out as part of the SKACH consortium through funding from SERI. The authors acknowledge the support of the Center for Scientific Computing (sciCORE) at the University of Basel, where part of these calculations were performed.